OpenStudy (anonymous):

f(m,n) = m^2 – n^2 is onto or not kindly tell me with detail because i don't understand how ?

OpenStudy (zzr0ck3r):

onto what?

OpenStudy (perl):

lets say the function is ZxZ -> Z

OpenStudy (perl):

f(m,n) is onto if for each element of y of Z there exists (m,n) of ZxZ such that f(m,n) = y

OpenStudy (perl):

so it is not onto if you can find some number such that f(m,n) is never equal to y . what about y = 2. is the difference of two squares ever equal to 2?

OpenStudy (perl):

also i am assuming, m,n are integers

OpenStudy (zzr0ck3r):

So in \(\mathbb{R}\) yes but not in \(\mathbb{Q}, \text{ or } \mathbb{Z} \)

OpenStudy (perl):

right, it depends on what the domain and codomain are.

OpenStudy (perl):

how do you type R in latex?

OpenStudy (perl):

without going into the equation editor

OpenStudy (zzr0ck3r):

normally at lower level we only talk about the real line, so if this is for something a little more advanced I am surprised the OP did not specify that set...

OpenStudy (zzr0ck3r):

`\(\mathbb{R}\)`

OpenStudy (perl):

\(\mathbb{R}\)

OpenStudy (zzr0ck3r):

:)

OpenStudy (perl):

how did you get that in grey?

OpenStudy (zzr0ck3r):

hmm I dont know if i can type it put a ` at the start of your code, and ` at the end

OpenStudy (zzr0ck3r):

hmm

OpenStudy (perl):

a\(\mathbb{R}\)a

OpenStudy (zzr0ck3r):

put the thing under ~ at the start of your code and the thing under ~ at the end

OpenStudy (perl):

zzrock, can you help me with my question, sorry i dont mean to be rude

OpenStudy (perl):

i am snatching you :)

OpenStudy (perl):

`\(\mathbb{R}\)`

OpenStudy (perl):

its basic algebra

OpenStudy (zzr0ck3r):

if I can, sure where is it?